OpenStudy (elleblythe):

How to find the exact value of cos(13pi/8)??

OpenStudy (anonymous):

half angle formula for this one

OpenStudy (anonymous):

\[\frac{13\pi}{8}\] is half of \(\frac{13\pi}{4}\)

OpenStudy (anonymous):

you good from there?

OpenStudy (elleblythe):

@satellite73 I was looking at the solution in my notes and it says it becomes sin(pi/8) at one point, how does that happen?